MiraAngel:

Help Asap! I am having so much trouble with this subject and I have multiple problems! Please help! lol

MiraAngel:

1 attachment
Aqual:

so you have 23 students, 20 of which wont hold office

MiraAngel:

yes

Aqual:

So take 23 and set that to your combinations of 1st office

Aqual:

Now, take that number down to 22 and that is how many can hold 2nd office * 23.

MiraAngel:

combinations as in ncr=n!/((n-r)! r!) ????

Aqual:

Huh?

Aqual:

I don't know the equation.

MiraAngel:

lol okay. I think I have an idea of what to do. Still a little lost tho. lol

Aqual:

You now have the combinations for 1st and 2nd office.

Aqual:

So 23 *22

Aqual:

Now you want to multiply that by 21 for how many students can hold 3rd office.

Aqual:

@Pixel

Aqual:

Is that right?

Aqual:

10626

Pixel:

i believe so

MiraAngel:

Okay! Thank you!

Kingtaker435:

ill help

Rigo:

ok

iphoneking13:

mmmmmmm

Shadow:

\[_{n}C _{k} = \frac{ n! }{ k! (n-k)! }\]

iphoneking13:

dam

Shadow:

Where, \[_{23}C _{3} = \frac{ 23! }{ 3! (20 - 3)! }\]

MiraAngel:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shadow Where, \[_{23}C _{3} = \frac{ 23! }{ 3! (20 - 3)! }\] \(\color{#0cbb34}{\text{End of Quote}}\) Thanks!

Shadow:

(Missed a 23) Where, \[_{23}C _{3} = \frac{ 23! }{ 3! (23 - 3)! }\]

Shadow:

Think my mind was already thinking 20 lol

MiraAngel:

lollll I got it. lolll

Shadow:

But basically you get: \[\frac{ 23! }{ 3! \times 20! } \rightarrow \frac{ 23 \times 22 \times 21 \times 20! }{ 3! \times 20! }\]

Shadow:

Becoming: \[\frac{ 23 \times 22 \times 21 }{ 3 \times 2 \times 1} = 1771\]

MiraAngel:

Thank you! I have another problem posted if you're able to help on it! :D